A manufacturer of electronics products is considering entering the telephone equipment business> It estimates that if it were to begin making wireless telephones, its short-run cost function would be as follows:

Q (thousands)

AVC ($)

AC ($)

MC ($)

13

37.90

45.59

31.90

14

37.60

44.74

33.70

15

37.50

44.17

36.10

16

37.60

43.85

39.10

17

37.90

43.78

42.70

18

38.40

43.96

46.90

19

39.10

44.36

51.70

20

40.00

45.00

57.10

 

1. Plot the average cost, average variable cost and marginal cost on a graph.

2. Suppose that the wholesale price of a wireless phone is currently $50, what is the profit maximizing output level?

3. How much is the total profit of this company?

4. Suppose that the firm does enter the market and that overtime increasing competition causes prices of telephone to fall to $35, how much is the new production level?

5. How much is the profit (loss) when the price drops to $35?

6. Should the firm stay in business or shut down in the short run if the price is $35?



1
Expert's answer
2021-12-15T22:35:31-0500






2.

P=50

MC=51.70

Q=19


3.

TR=19×50=950

AC"=44.36"

"44.36=\frac{TC}{19}"

"TC=842.84"

"Profit=TR-TC"

"=950-842.84=107.26"


4.

P=MC

"P=35"

Q=15


5.

"TR=15\times35=525"

"AC=44.17"

"44.17=\frac{TC}{15}"

"TC=662.55"

Profit(Loss)"=525-662.55"

"=-137.55"


6.

In the short run, the firm should shut down because average variable costs exceed the price.